OpenStudy (anonymous):

2x-1 is a factor of f(x) =12x^3+kx^3+mx-3. as well when f(x) is divided by x+2 the remainder is -65. determine the value of m and k.

OpenStudy (anonymous):

if 2x-1 is a factor, then 2x-1 = 0 => x = 1/2 use synthetic division to find an equation in k and m. then if x+2 gives a remainder of -65 => when x = -2, f(-2) = -65. you can use synthetic division again to get another equation in k and m. now you have a system and can solve for k and m.

OpenStudy (dumbcow):

use synthetic division: 2x-1 ---> x = 1/2 , remainder = 0 x+2 --> x = -2 , remainder = -65 1/2 | 12 k m -3 6 1/2(k+6) m/2 +(k+6)/4 ----------------------------------------------------- 12 k+6 m+(k+6)/2 0 ---> m/2 + (k+6)/4 = 3 ----> m = (6-k)/2 -2 | 12 k m -3 -24 48-2k 4k -2m -96 ------------------------------------------------ 12 k-24 m+48-2k -65 -----> 4k -2m -99 = -65 ---> 2k - m = 17 Substitute: 2k - (6-k)/2 = 17 5k - 6 = 34 k = 8 ----> m = -1

OpenStudy (anonymous):

wouldn't we have to divide m/2+(k+6)/4 by 2 because we just divided by 1/2 and not 2

OpenStudy (anonymous):

we divided by x-1/2 not 2

OpenStudy (anonymous):

well, i've never learnt synthetic division as its out of my syllabus but there's a simpler method to simply sub in the x values as 2x-1 is a factor, 2x=1 x=1/2 you got that, yes. so simply make replace all the x in your equation as 1/2 factor means your equation will be = 0 therefore 12(1/2)^3+k(1/2)^3+m(1/2)-3=0 simplify this first now, for two unknowns, you need two equations do the same thing for x+2, where x is now = -2 and instead of 0, use -65 substitute and solve

OpenStudy (dumbcow):

sorry i dont follow....m/2 +(k+6)/4 is just the result of using synthetic division

OpenStudy (anonymous):

but we are divding the polynomial by x+1/2 which comes out of 2(x+1/2), so don't we also divide by 2 that is factored out?

OpenStudy (dumbcow):

yeah @Auxuris that is prob easier

OpenStudy (anonymous):

yeah thanks

OpenStudy (dumbcow):

ahh i see, well thankfully remainder is 0 so dividing by 2 does not affect the remainder at all

OpenStudy (anonymous):

but if the remainder was not zero we would have to divide by 2 right?

OpenStudy (dumbcow):

yes but only the remainder would be divided by 2..... not the m/2 +(k+6)/4 part the synthetic div process remains same

OpenStudy (anonymous):

so if m/2+(k+6)/4 + (-3) = 5, we would divide 5/2?

OpenStudy (dumbcow):

yes so you would equal it to 5/2 instead of 5

OpenStudy (anonymous):

okay, makes sense

OpenStudy (anonymous):

thanks for your help!

OpenStudy (dumbcow):

yw